OpenStudy (anonymous):

Ben made some home improvements for $9,375 using a six-month deferred payment plan with an interest rate of 19.42%. What is the balance after the deferment period if no payments are made? $10,322.95 $9,375.00 $11,195.63 $10,244.95 I know it's B = P (1 + i)^nt... but I can't get it right!!

OpenStudy (anonymous):

try using 0.5 as exponent

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

A=P(1+r)^t A=9375(1.1942)^0.5

OpenStudy (anonymous):

I think that works! :D THANKS!

OpenStudy (anonymous):

thanks for the medal :D

OpenStudy (anonymous):

Welcome... :)